Required Information [The following information applies to the questions displayed below.] Part 1 of 4 1.64 points Henrich isb. His $452,000 of taxable income includes $2.000 of long-term capital gain that is taxed at preferential rates. Income tax Nc. His $452,000 of taxable income includes $43,000 of long-term capital gain that is taxed at preferential rates. Income taxd. Henrich has $196,000 of taxable income, which includes $50,400 of long-term capital gain that is taxed at preferential ratEstates and Trusts If taxable income is The tax is: But not over: over: $ 0 $ 2,550 10% of taxable income $ 2,550 $ 9,150 $25Tax Rates for Net Capital Gains and Qualified Dividends Rate* | 0% Married Filing Jointly SO - $77,200 $77,201 - $479,000 $47INUIviuudis Schedule X-Single If taxable income is over: But not over: The tax is: S 0 S 9,525 10% of taxable income $ 9,525
Answer #1

calculation of Henrich tax liability is different scenarios

a) in this scenario all income is coming from employer as salary and there is no capital gain so no net investment tax

Calculation of federal Tax liability

Particulars

Amount($)

Gross salary

452000

Income before exemption and deduction

452000

less personal exemption

-0

all personsal exemption are disallowed from year 2018

less Standard  deduction

-12000

standard deduction allowed $ 12000 for single tax filler

Taxable income

440000

(452000-12000)

Taxes on Taxable income

Tax liability on base bracket

45689.50

taken from the table

Tax liability on excess over base@35%

84000

($440000-$200000 )*35%

Total federal income tax liability

129689.5

45689.50+84000=129689.5
Net investment Income tax Liability 0

B) in this case 452000 income includes 2000 long term gain i.e is calculated at preferential rates

Long term gain are gain which arises on assets or common stock held for more than year

Long term gain on  $ 2000 in case of Mr henrich will be @ 20% as his taxable income is greater than $425800

As long term Gain is arrising from estate and trust so as per federal tax rules it can be taxed in either in hands of henrich or estate which have lower tax on capital gain

as estate total income is $ 2000 ,so tax liability on its long term capital gain is 0%

so its better the long capital gain should be taxed in hand of estate and trust

calculation of federal income tax and Net Investment tax Liability

Calculation of federal Tax liability

Particulars

Amount($)

Gross salary

450000

Income before exemption and deduction

450000

less personal exemption

-0

less standard  deduction

-12000

standard deduction allowed $ 12000 for single tax filler

Taxable income before Lt gain And Dividends

438000

450000-12000

Taxes on Taxable income before LT Gain and dividends :

Tax liability on base bracket

45689.50

Tax liability on excess over base@35%

83300

$(438000-200000)*35%

Tax on Taxable income before LT Gain and dividends

128990

45689.50+83300=128989.5

Taxes on LT Gain and dividends :

Lt gain

2000

dividends

0

2000

Tax on LT Gain and dividends :@0%

0

as long term gain is received from estate and trust and it is taxable in hands of estate and trust it will be exempt in hands of henrich

total federal tax liability

128990

=128990+0

net investment tax liability 0 as net investment should be 200000 or AGI should be more than 210000

c) In this case 452000 taxable income includes 43000 of long term gain

Gross salary = 452000- 43000

= 409000

Calculation of federal Tax liability

Particulars

Amount($)

Gross salary

409000

452000-43000

Income before exemption and deduction

409000

less personal exemption

-0

less standard  deduction

-12000

standard deduction allowed $ 12000 for single tax filler

Taxable income before Lt gain And Dividends

397000

409000-12000

Taxes on Taxable income before LT Gain and dividends :

Tax liability on base bracket

45689.50

Tax liability on excess over base@35%

68950

$(397000-200000)*35%

Tax on Taxable income before LT Gain and dividends

114640

45689.50+68950=114639.5

Taxes on LT Gain and dividends :

Lt gain

43000

dividends

0

43000

Tax on LT Gain and dividends :@20 %

8600

as long term gain is received from estate and trust and it is taxable in hands henrich as tax liability in hands of estate for long term gain is more than henrich

total federal tax liability

123240

=114640+8600

net investment tax liability 0 as net investment should be 200000 or AGI should be more than 210000

d) 196000 which includes 50400 long ter gain

Calculation of federal Tax liability

Particulars

Amount($)

Gross salary

145600

196000-50400
adjusted Gross income 210000

Income before exemption and deduction

210000

less personal exemption

-0

less standard  deduction

-12000

standard deduction allowed $ 12000 for single tax filler

Taxable income before Lt gain And Dividends

198000

210000-12000

Taxes on Taxable income before LT Gain and dividends :

Tax liability on base bracket

32089.50

Tax liability on excess over base@32%

12960

$(198000-157500)*32%

Tax on Taxable income before LT Gain and dividends

45050

32089.50+12960=45049.5

Taxes on LT Gain and dividends :

Lt gain

50400

dividends

0

50400

Tax on LT Gain and dividends :@15 %

7560

as long term gain is received from estate and trust and it is taxable in hands henrich as tax liability in hands of estate for long term gain is more than henrich

total federal tax liability

52610

45050+7560

net investment tax liability 380

as net investment should be 200000 or AGI should be more than 210000 so he is liable to pay 3.8 % tax

AGI EXCESS INVESTMENT =210000-200000=10000

CAPITAL GAIN =50400

THE LOWER VALUE $10000 IS TAXABLE @3.8%